### v1

@@ 2.1 System Design
(a) Leader module (b) Follower module
@@
Figure 2: Experimental platform used to evaluate the proposed NIMA framework. (a) The leader module consists of a dual-haptic console (Omega.7, Force Dimension) used by the surgeon to teleoperate the robotic arms through motion commands, with real-time feedback displayed on a monitor. (b) The follower module comprises two Kinova Gen3 robotic arms equipped with 6-DoF Bota force–torque sensors, custom instrument adapters, and a translucent mannequin containing a soft-tissue surrogate. The setup replicates a realistic minimally invasive surgical environment for evaluating tool–tissue interaction forces, motion tracking, and haptic feedback performance.
@@
The experimental apparatus was designed to simulate a realistic surgical environment by integrating advanced technologies and components, as depicted in Fig. 2. The setup included a custom mannequin embedded with tissue-representative objects, specially designed surgical tools, two Kinova Gen3 robotic arms (7 DOF), dedicated carts for robotic arm stability, two Omega.7 haptic controllers (Force Dimension), three 6-axis force–torque sensors (SensONE, Bota Systems), and an adaptable, in-house-designed surgical console. The Kinova Gen3 robotic arms were essential for delivering precise and controlled movements of the surgical instruments, replicating the dexterity of human hands to ensure an authentic experimental environment for surgical tasks. Specialized carts were engineered to support the robotic arms, providing both stability and flexibility in positioning the system during experiments. The surgical tools were custom-designed to integrate seamlessly with the robotic arms while adhering to clinical standards and were powered by four Dynamixel actuators (ROBOTIS) to achieve fine-grained control and precision during operation. The Omega.7 haptic controllers enabled the operator to intuitively interact with the robotic arms, translating hand movements into precise surgical motions. Their high positional accuracy played a crucial role in creating realistic and user-friendly surgical simulations. A translucent mannequin, as shown in Fig. 2a, incorporated tissue-representative objects to replicate the structural complexities of real surgical scenarios, including pick-and-place and suturing tasks, creating an environment that closely mirrors actual surgical challenges. Force data from the simulated tasks were captured using Bota force sensors mounted on the tips of the robotic arms. These sensors provided valuable insights into the mechanical interactions between the surgical tools and the tissue-representative objects. The custom-designed surgical console, shown in Fig. 2b, acted as the central control unit, seamlessly integrating the robotic arms, haptic controllers, surgical tools, Dynamixel actuators, and force sensors, ensuring efficient communication and coordination among all components.
@@
Figure 3: System architecture of the proposed NIMA-enabled robot-assisted laparoscopy setup. The leader module (left) includes dual haptic devices (Omega.7) that transmit motion commands and receive force feedback estimated by the NIMA. The follower module (right) consists of two robotic arms equipped with 6-DoF force–torque sensors, an endoscope, and optical markers tracked by an NDI system for spatial registration. Force and motion data are processed through a NN for tool-tip force extraction and used by NIMA to render stable, high-fidelity haptic.
@@
The system architecture, illustrated in Fig. 3, enabled coordinated control by transmitting motion commands from the haptic devices to both the robot and instrument controllers. The robotic arms executed position commands, while the Dynamixel actuators controlled the orientation of the surgical tool tip, with each task managed by a dedicated control module. A PID controller was implemented for precise position and velocity regulation, ensuring accurate tool-tip alignment during operation. The robot controller consisted of two main components: one for hand–eye coordination and another for robot communication. This structure synchronized the robotic arms, optical tracker, endoscope, and haptic devices with the surgeon’s visual perspective to maintain cohesive and intuitive teleoperation. Bota 6-DoF force–torque sensors, mounted between the robots’ end effectors and the custom instrument adapters, measured the interaction forces between the surgical tools and tissue, as well as the frictional forces generated at the Remote Center of Motion (RCM). To maintain a stable entry point on the patient’s body, translational movements of the surgeon’s hand were converted into rotational motions about the RCM. The friction forces induced during these translational motions were calculated and subtracted from the total measured forces to isolate the true tool-tip forces. A NN module was then developed to estimate these tool-tip forces based on the signals acquired from the force sensors at the robot end effectors, as described in Section 2.2. The resulting tool–tissue interaction forces were transmitted to the NIMA framework, which generated adaptive haptic feedback to the surgeon’s hands, allowing the feedback intensity to be adjusted according to user preference.
@@
For incorporating force feedback, accurately identifying the force sensor system is essential for applying gravitational biasing, which ensures precise robotic instrument movements. This study employs system identification using IMU data to determine the spatial orientation of the end effector, enabling effective gravitational biasing methods that counteract the effects of gravity on the robotic tools. To achieve this, an Integral-Free Spatial Orientation Estimation Method is introduced, which utilizes IMU sensors to calculate the end effector’s orientation in real time without relying on data integration from other sources. Using this IMU data, the system effectively calculates the end effector’s roll, pitch, and yaw, critical for compensating gravitational forces acting on the robotic instruments. For simplicity, this method computes roll and pitch angles using accelerations in three directions, avoiding the integration of angular velocity over time 9397346. Real-time angles are calculated as:
@@ 2.2.1 Coordinate Systems Correspondence
$^{\{\textbf{S}_{1}\}}\textbf{R}_{\{\textbf{S}_{2}\}}=^{\{\textbf{S}_{1}\}}\textbf{R}_{\{\textbf{KBF}\}}{\penalty 10000\ }^{\{\textbf{KBF}\}}\textbf{R}_{\{M\}}{\penalty 10000\ }^{\{\textbf{M}\}}\textbf{R}_{\{\textbf{C}\}}{\penalty 10000\ }^{\{\textbf{C}\}}\textbf{R}_{\{\textbf{S}_{2}\}}$ (11)
@@ 2.2.2 Neural Tool-Tissue Force Estimation
where $\textbf{F}_{1}\in\mathbb{R}^{1\times 3}$, $\textbf{F}_{2}\in\mathbb{R}^{1\times 3}$, and $\textbf{F}_{3}\in\mathbb{R}^{1\times 3}$ are the forces at the robotic arm, RCM point, and tip forces. During the experiment, $\textbf{F}_{1}$$\textbf{F}_{3}$, and quaternions of the robotic arm q were captured. These parameters create a dataset of $30000\times 10$ for a fully connected feedforward NN designed for this regression task where $(\textbf{F}_{1})_{3\times 30000}$, $(\textbf{q})_{4\times 30000}$ with 7 features and 30000 observations were the inputs and $(\textbf{F}_{3})_{3\times 30000}$ with 3 features and 30000 observations were the output.
Figure 5: Internal view of the mannequin used in Experiment 2, illustrating the setup for acquiring tool–tissue interaction data to train the NN model. A surgical instrument, inserted through the upper access port, interacts with a flexible tissue representative mounted on a 6-DoF Bota force–torque sensor. An endoscope provides visual feedback to monitor tool motion and contact inside the surgical workspace.
@@ 2.3 NIMA-based Force Reconstruction
where $\textbf{m}_{i}\in\mathbb{R}^{1\times 3N}$ is a vector of impedance parameters and $\textbf{0}\in\mathbb{R}^{1\times 3N}$ is a zero vector. This model assumes no cross-talk between the orthogonal tool-tissue forces along $x$, $y$, and $z$ axes. Cross-talk effects, if needed, can be modeled by including additional non-zero elements in M. To compute M in real-time, a rolling time window of $\delta t=300$ ms was adopted. Delays under 300 ms are generally imperceptible to surgeons in leader-follower setups Perez2007, making this an effective worst-case scenario for updating impedance parameters. The NIMA parameters M are continuously updated using a rolling dataset of $n$ sample forces $\hat{\textbf{F}}\in\mathbb{R}^{3\times n}$ and motion commands $\hat{\textbf{X}}$:
@@
The peg transfer task, a core element of the Fundamentals of Laparoscopic Surgery (FLS) curriculum Fried2007, was chosen for evaluation. This task involved capturing the robotic arms’ positions, orientations, and velocities using two Omega.7 haptic controllers, synchronized to the operator’s motion commands at a refresh rate of 1 kHz with a programmed delay of 300 ms. Force feedback, essential for realistic haptic interaction, was recorded via force sensors (SensOne, Bota Systems) at 2 kHz. These data were used to identify NIMA parameters and apply Eq. 16 to compute and render the desired 3D force, $\textbf{f}_{d}$, on the haptic device. To optimize the polynomial degree $N$ for NIMA, five parallel threads with $N=1\cdots 5$ were executed, with the optimal $N$ for each time window selected based on minimizing the 3D force reconstruction error.
@@
Figure 7: Performance of the trained NN model in predicting tool-tip forces after removing frictional effects at the RCM. The plots show measured forces (ground truth) and model predictions along the (a) $X$, (b) $Y$, and (c) $Z$ directions, with shaded regions representing the 95% confidence intervals. The close overlap between measured and predicted forces demonstrates the NN model’s high accuracy and consistency across all axes, validating its capability to capture nonlinear tool–tissue interaction dynamics.
3 Validation Studies
@@ 3.3 Validation III: NIMA-based Force Reconstruction
The effectiveness of our NIMA in accurately rendering forces on haptic devices is illustrated in Fig. 10. Figure 10 compares the forces rendered by the haptic device to the actual forces measured by force sensors. The implementation of NIMA achieved a MAE of 0.01 N, demonstrating high fidelity in force feedback. The errors followed a normal distribution with a SD of 0.02 N. This performance significantly exceeds that of the Linear IMA, which recorded an MAE of 0.2 N and an SD of 0.4 N. A comparative assessment reveals an 95% improvement in accuracy with NIMA. This enhancement underscores the superior precision of the nonlinear approach in generating force feedback, setting a new benchmark for realism and immersion in haptic interactions. The substantial reduction in MAE highlights the effectiveness of integrating nonlinear dynamics into impedance matching processes, improving both the quality and reliability of haptic feedback. Furthermore, Fig. 10 demonstrates how NIMA addresses the kick-back behavior by rendering no forces for movements with velocities below 1 mm/s. Consequently, when the user releases the haptic device, no force is applied to the handles, ensuring they remain stationary. However, resisting forces are rendered with high accuracy once user input exceeds the velocity threshold, preserving NIMA’s precision. A detailed analysis of the algorithm’s performance throughout the experiment revealed a noteworthy preference for nonlinear models, with a nonlinear fit being selected in over 64% of the evaluated time windows. This reliance on nonlinear approaches highlights the limitations of the Linear IMA model in capturing the intricate, dynamic interactions between surgical tools and tissue. The complexity and variability of these interactions exceed the representational capabilities of linear models, emphasizing the nuanced nature of tool-tissue forces encountered during laparoscopic procedures. These findings suggest that temporal variations in these forces, which are crucial for realistic haptic feedback, are better captured through nonlinear modeling, enabling a more accurate and responsive simulation of surgical scenarios.
@@
The experimental findings confirm that the proposed NIMA significantly enhances the stability, transparency, and realism of force feedback in RAMIS. A particularly noteworthy outcome across all trials was the complete absence of the phenomenon commonly known as the haptic kick, a sudden and undesirable jerk experienced by the operator when releasing the haptic interface. This behavior, often observed in conventional force feedback systems, typically results from the delayed response of the controller or the residual energy stored within the actuator–sensor loop. In the case of NIMA, the elimination of this artifact stems from the rapid convergence of the system’s output force vector X to zero within a short time interval $\delta t$ once the operator releases the haptic device. Consequently, the desired force $\textbf{f}^{d}$ also decays smoothly to zero, ensuring that no reactive forces are transmitted back to the user. This dynamic behavior guarantees passive stability and prevents unintended force reflections, thereby improving both operator comfort and system safety. These results are consistent with the impedance-based framework previously described by Sayadi et al. sayadi2020impedance, but the nonlinear formulation implemented in NIMA allows a more robust and adaptive response across a broader range of tool–tissue interaction conditions. Beyond mitigating haptic kickback, NIMA demonstrates strong adaptability and precision in representing nonlinear contact dynamics between the surgical instrument and the tissue phantom. Traditional linear impedance control approaches often oversimplify these interactions, limiting their ability to reproduce complex soft-tissue behaviors and variable stiffness profiles encountered during real surgical procedures. By incorporating higher-order nonlinear terms and updating impedance parameters in real time, NIMA effectively maintains consistent performance under varying operational and material conditions. The enhanced transparency of the system allows the operator to perceive genuine tool–tissue forces without interference from control-induced artifacts, improving tactile perception during teleoperation. This capability is crucial in delicate surgical manipulations, where even small unintentional force fluctuations can lead to tissue deformation or procedural errors. The robustness and stability demonstrated by NIMA across multiple experimental conditions suggest its strong potential for integration into clinical robotic systems and surgical training simulators. By delivering stable, high-fidelity, and realistic haptic feedback, NIMA moves one step closer toward achieving the level of tactile awareness required for safe, precise, and intuitive human–robot interaction in next-generation RAMIS platforms.

### v2

@@ 2.1 System Design
(a) Follower module (b) Leader module
@@
Figure 2: Experimental platform used to evaluate the proposed NIMA framework. (a) The follower module comprises two Kinova Gen3 robotic arms equipped with 6-DoF Bota force–torque sensors, custom instrument adapters, and a translucent mannequin containing a soft-tissue surrogate. (b) The leader module consists of a dual-haptic console (Omega.7, Force Dimension) used by the surgeon to teleoperate the robotic arms through motion commands, with real-time feedback displayed on a monitor. The setup replicates a realistic minimally invasive surgical environment for evaluating tool–tissue interaction forces, motion tracking, and haptic feedback performance.
@@
The experimental apparatus was designed to simulate a realistic surgical environment by integrating advanced technologies and components, as depicted in Fig. 2. The setup included a custom mannequin embedded with tissue–representative objects, specially designed surgical tools, two Kinova Gen3 robotic arms (7 DOF), dedicated carts for robotic arm stability, two Omega.7 haptic controllers (Force Dimension), three six–axis force–torque sensors (SensONE, Bota Systems), and an adaptable, in-house-designed surgical console. The Kinova Gen3 robotic arms were essential for delivering precise and controlled movements of the surgical instruments, replicating the dexterity of human hands to ensure an authentic experimental environment for surgical tasks. Specialized carts were engineered to support the robotic arms, providing both stability and flexibility in positioning the system during experiments. The surgical tools were custom-designed to integrate seamlessly with the robotic arms while adhering to clinical standards and were powered by four Dynamixel actuators (ROBOTIS) to achieve fine-grained control and precision during operation. The Omega.7 haptic controllers enabled the operator to intuitively interact with the robotic arms, translating hand movements into precise surgical motions. Their high positional accuracy played a crucial role in creating realistic and user-friendly surgical simulations. A translucent mannequin, as shown in Fig. 2a, incorporated tissue-representative objects to replicate the structural complexities of real surgical scenarios, including pick and place and suturing tasks, creating an environment that closely mirrors actual surgical challenges. Force data from the simulated tasks were captured using Bota force sensors mounted on the tips of the robotic arms. These sensors provided valuable insights into the mechanical interactions between the surgical tools and the tissue-representative objects. The custom-designed surgical console, shown in Fig. 2b, acted as the central control unit, seamlessly integrating the robotic arms, haptic controllers, surgical tools, Dynamixel actuators, and force sensors, ensuring efficient communication and coordination among all components.
@@
Figure 3: System architecture of the proposed NIMA-enabled robot-assisted laparoscopy setup. The leader module (left) includes dual haptic devices (Omega.7) that transmit motion commands and receive force feedback estimated by the NIMA. The follower module (right) consists of two robotic arms equipped with 6-DoF force-torque sensors, an endoscope, and optical markers tracked by an NDI system for spatial registration. Force and motion data are processed through a NN for tool-tip force extraction and used by NIMA to render stable, high-fidelity haptic.
@@
The system architecture, illustrated in Fig. 3, enabled coordinated control by transmitting motion commands from the haptic devices to both the robot and instrument controllers. The robotic arms executed position commands, while the Dynamixel actuators controlled the orientation of the surgical tool tip, with each task managed by a dedicated control module. A PID controller was implemented for precise position and velocity regulation, ensuring accurate tool-tip alignment during operation. The robot controller consisted of two main components: one for hand–eye coordination and another for robot communication. This structure synchronized the robotic arms, optical tracker, endoscope, and haptic devices with the surgeon’s visual perspective to maintain cohesive and intuitive teleoperation. Bota 6-DoF force–torque sensors, mounted between the robots’ end effectors and the custom instrument adapters, measured the interaction forces between the surgical tools and tissue, as well as the frictional forces generated at the RCM. To maintain a stable entry point on the patient’s body, translational movements of the surgeon’s hand were converted into rotational motions about the RCM. The friction forces induced during these translational motions were calculated and subtracted from the total measured forces to isolate the true tool-tip forces. A NN module was then developed to estimate these tool-tip forces based on the signals acquired from the force sensors at the robot end effectors, as described in Section 2.2. The resulting tool–tissue interaction forces were transmitted to the NIMA framework, which generated adaptive haptic feedback to the surgeon’s hands, allowing the feedback intensity to be adjusted according to user preference.
@@
For incorporating force feedback, accurately identifying the force sensor system is essential for applying gravitational biasing, which ensures precise robotic instrument movements. This study employs system identification using IMU data to determine the spatial orientation of the end effector, enabling effective gravitational biasing methods that counteract the effects of gravity on the robotic tools. To achieve this, an Integral-Free Spatial Orientation Estimation Method is introduced, which utilizes IMU sensors to calculate the end effector’s orientation in real time without relying on data integration from other sources. Using this IMU data, the system effectively calculates the end effector’s roll, pitch, and yaw, critical for compensating gravitational forces acting on the robotic instruments. For simplicity, this method computes roll and pitch angles using accelerations in three directions, avoiding the integration of angular velocity over time [19]. Real-time angles are calculated as:
@@ 2.2.1 Coordinate Systems Correspondence
$^{\{\textbf{S}_{1}\}}\textbf{R}_{\{\textbf{S}_{2}\}}=^{\{\textbf{S}_{1}\}}\textbf{R}_{\{\textbf{KBF}\}}{~}^{\{\textbf{KBF}\}}\textbf{R}_{\{M\}}{~}^{\{\textbf{M}\}}\textbf{R}_{\{\textbf{C}\}}{~}^{\{\textbf{C}\}}\textbf{R}_{\{\textbf{S}_{2}\}}$ (11)
@@ 2.2.2 Neural Tool-Tissue Force Estimation
where $\textbf{F}_{1}\in\mathbb{R}^{1\times 3}$, $\textbf{F}_{2}\in\mathbb{R}^{1\times 3}$, and $\textbf{F}_{3}\in\mathbb{R}^{1\times 3}$ are the forces at the robotic arm, RCM point, and tip forces. During the experiment, $\textbf{F}_{1}$, $\textbf{F}_{3}$, and quaternions of the robotic arm q were captured. These parameters create a dataset of $30000\times 10$ for a fully connected feedforward NN designed for this regression task where $(\textbf{F}_{1})_{3\times 30000}$, $(\textbf{q})_{4\times 30000}$ with 7 features and 30000 observations were the inputs and $(\textbf{F}_{3})_{3\times 30000}$ with 3 features and 30000 observations were the output.
Figure 5: Internal view of the mannequin used in Experiment 2, illustrating the setup for acquiring tool-tissue interaction data to train the NN model. A surgical instrument, inserted through the upper access port, interacts with a flexible tissue representative mounted on a 6-DoF Bota force–torque sensor. An endoscope provides visual feedback to monitor tool motion and contact inside the surgical workspace.
@@ 2.3 NIMA-based Force Reconstruction
where $\textbf{m}_{i}\in\mathbb{R}^{1\times 3N}$ is a vector of impedance parameters and $\textbf{0}\in\mathbb{R}^{1\times 3N}$ is a zero vector. This model assumes no cross-talk between the orthogonal tool-tissue forces along $x$, $y$, and $z$ axes. Cross-talk effects, if needed, can be modeled by including additional non-zero elements in M. To compute M in real-time, a rolling time window of $\delta t=300$ ms was adopted. Delays under 300 ms are generally imperceptible to surgeons in leader-follower setups [32], making this an effective worst-case scenario for updating impedance parameters. The NIMA parameters M are continuously updated using a rolling dataset of $n$ sample forces $\hat{\textbf{F}}\in\mathbb{R}^{3\times n}$ and motion commands $\hat{\textbf{X}}$:
@@
The peg transfer task, a core element of the Fundamentals of Laparoscopic Surgery (FLS) curriculum [12], was chosen for evaluation. This task involved capturing the robotic arms’ positions, orientations, and velocities using two Omega.7 haptic controllers, synchronized to the operator’s motion commands at a refresh rate of 1 kHz with a programmed delay of 300 ms. Force feedback, essential for realistic haptic interaction, was recorded via force sensors (SensOne, Bota Systems) at 2 kHz. These data were used to identify NIMA parameters and apply Eq. 16 to compute and render the desired 3D force, $\textbf{f}_{d}$, on the haptic device. To optimize the polynomial degree $N$ for NIMA, five parallel threads with $N=1\cdots 5$ were executed, with the optimal $N$ for each time window selected based on minimizing the 3D force reconstruction error. This ensured accurate force feedback and enhanced the surgical simulation’s realism.
@@
Figure 7: Performance of the trained NN model in predicting tool-tip forces after removing frictional effects at the RCM. The plots show measured forces (ground truth) and model predictions along the (a) $X$, (b) $Y$, and (c) $Z$ directions, with shaded regions representing the 95% confidence intervals. The close overlap between measured and predicted forces demonstrates the NN model’s high accuracy and consistency across all axes, validating its capability to capture nonlinear tool-tissue interaction dynamics.
3 Results
@@ 3.3 Validation III: NIMA-based Force Reconstruction
The effectiveness of our NIMA in accurately rendering forces on haptic devices is illustrated in Fig. 10, which compares the forces rendered by the haptic device to the actual forces measured by force sensors. The implementation of NIMA achieved a MAE of 0.01 N, demonstrating high fidelity in force feedback. The errors followed a normal distribution with a SD of 0.02 N. This performance significantly exceeds that of the Linear IMA [37], which recorded an MAE of 0.2 N and an SD of 0.4 N. A comparative assessment reveals a 95% improvement in accuracy with NIMA. This enhancement underscores the superior precision of the nonlinear approach in generating force feedback, setting a new benchmark for realism and immersion in haptic interactions. The substantial reduction in MAE highlights the effectiveness of integrating nonlinear dynamics into impedance matching processes, improving both the quality and reliability of haptic feedback. Furthermore, Fig. 10 demonstrates how NIMA addresses the kickback behavior by rendering no forces for movements with velocities below 1 mm/s. Consequently, when the user releases the haptic device, no force is applied to the handles, ensuring they remain stationary. However, resisting forces are rendered with high accuracy once user input exceeds the velocity threshold, preserving NIMA’s precision. A detailed analysis of the algorithm’s performance throughout the experiment revealed a noteworthy preference for nonlinear models, with a nonlinear fit being selected in over 64% of the evaluated time windows. This reliance on nonlinear approaches highlights the limitations of linear impedance models in capturing the intricate, dynamic interactions between surgical tools and tissue. The complexity and variability of these interactions exceed the representational capabilities of linear models, emphasizing the nuanced nature of tool-tissue forces encountered during laparoscopic procedures. These findings suggest that temporal variations in these forces, which are crucial for realistic haptic feedback, are better captured through nonlinear modeling, enabling a more accurate and responsive simulation of surgical scenarios.
4 Discussion
@@
The experimental findings confirm that the proposed NIMA significantly enhances the stability, transparency, and realism of force feedback in RAMIS. A particularly noteworthy outcome across all trials was the complete absence of the phenomenon commonly known as the haptic kick, a sudden and undesirable jerk experienced by the operator when releasing the haptic interface. This behavior, often observed in conventional force feedback systems, typically results from the delayed response of the controller or the residual energy stored within the actuator–sensor loop. In the case of NIMA, the elimination of this artifact stems from the rapid convergence of the system’s output force vector X to zero within a short time interval $\delta t$ once the operator releases the haptic device. Consequently, the desired force $\textbf{f}^{d}$ also decays smoothly to zero, ensuring that no reactive forces are transmitted back to the user. This dynamic behavior guarantees passive stability and prevents unintended force reflections, thereby improving both operator comfort and system safety. These results are consistent with the impedance-based framework previously described by Sayadi et al. [37], but the nonlinear formulation implemented in NIMA allows a more robust and adaptive response across a broader range of tool–tissue interaction conditions. Beyond mitigating haptic kickback, NIMA demonstrates strong adaptability and precision in representing nonlinear contact dynamics between the surgical instrument and the tissue phantom. Traditional linear impedance control approaches often oversimplify these interactions, limiting their ability to reproduce complex soft-tissue behaviors and variable stiffness profiles encountered during real surgical procedures. By incorporating higher–order nonlinear terms and updating impedance parameters in real time, NIMA effectively maintains consistent performance under varying operational and material conditions. The enhanced transparency of the system allows the operator to perceive genuine tool–tissue forces without interference from control-induced artifacts, improving tactile perception during teleoperation. This capability is crucial in delicate surgical manipulations, where even small unintentional force fluctuations can lead to tissue deformation or procedural errors. The robustness and stability demonstrated by NIMA across multiple experimental conditions suggest its strong potential for integration into clinical robotic systems and surgical training simulators. By delivering stable, high-fidelity, and realistic haptic feedback, NIMA moves one step closer toward achieving the level of tactile awareness required for safe, precise, and intuitive human–robot interaction in next-generation RAMIS platforms.